A baseball is hit with an initial speed of 35 m/s at an angle of 37 degrees with respect to the horizontal. How long does it take to hit the ground. You may assume that the batter hits the ball 1 meter above the ground
Answer #1

a)


for vertical motion,using the equation of motion,

s = u*t+0.5*a*t^2

where u = initial (vertical)velocity = 35*sin(37 degrees)

s = v ertical displacement = -1 m <----- ve sign for downward motion

a = -9.8 m/s2


So, -1 = 35*sin(37 degrees)*t + 0.5*(-9.8)*t^2

So, t = 4.35 s <-------------answer
Answer #2

Vo = 35m/s @ 37 deg.

Yo = ver. = 35sin37 = 21.06m/s.

h = (0 - (21.06)^2) / -19.6 = 22.63m
Answer #3

First it will move in a projectile path,


Vertical Velocity = Vsina


Time of Flight (T1)= 2Vsina/g


after that it will just fall freely for a height of 1m, Time (T2) = (2/g)^1/2


Therefore total time = T1 + T2 = 2*35*sin37/(9.8) + (2/9.8)^1/2

= 4.75 sec